You are given the following balance sheet of the Summer Bank Balance sheet of the Summer...

  1. You are given the following balance sheet of the Summer Bank

Balance sheet of the Summer bank

Assets

Liabilities

Cash $ 8,000


Deposited with the Fed $ 6,000

Loans $116,000

Deposits $ 80,000

Capital $ 50,000

Total $130,000

Total $ 130,000

The required reserve ratio (RRR) on all deposits is 10%

a. What, if any, are this bank's excess reserves?

b. How much new amount of loan will this bank be able to create because of the excess reserves?

c. How much new amount of loan the entire banking system be able to create because of the excess reserves?

d. What would be the excess reserves of this bank after the RRR is changed to 8%?

e. How much new amount of loan will this bank be able to create with the RRR of 8%?

f. How much new amount of loan the entire banking system be able to create because of the excess reserves?
Answer #1

Answer :-

a). According to the table ,

Deposits = $ 80,000

Required reserve ratio which is the percentage of deposit that must hold by banks = 10% = 0.10

Then ,
required reserve = Deposits × Reserve ratio

Required Reserve = 80000 × 0.10= 8000

Required Reserve = $ 8000


Actual Deposit = Cash + Deposit in Fed
Actual Deposit = $8000 + $6000 = $14000

Excess Reserve

= Actual Deposits - Required Reserve

Excess Reserve = 14000 - 8000 =

Excess Reserve = $ 6000

b) $ 6000 amount of loan will this bank be able to create because of the excess reserves.


c) The maximum amount the banking system might lend =
Excess reserve × 1/ Required reserve ratio

= 6000× 1/0.25
= 6000× 4
= $ 24000

The maximum amount the banking system might create = $24,000


d). According to the table ,

Deposits = $ 80,000

Required reserve ratio which is the percentage of deposit that must hold by banks = 8% = 0.08

Then ,
required reserve = Deposits × Reserve ratio

Required Reserve = 80000 × 0.08= 6400

Required Reserve = $ 6400


Actual Deposit = Cash + Deposit in Fed
Actual Deposit = $8000 + $6000 = $14000

Excess Reserve

= Actual Deposits - Required Reserve

Excess Reserve = 14000 - 6400

Excess Reserve = $ 7600

e) $ 7600 amount of loan will this bank be able to create because of the excess reserves.


f) The maximum amount the banking system might lend =
Excess reserve × 1/ Required reserve ratio

= 7600× 1/0.25
= 7600× 4
= $ 30,400

The maximum amount the banking system might create = $ 30,400